Please help me with this problem.
6m/7+1/2=1/2

Answer by jai_kos(139)

6m/7 + 1/2 = 1/2
Subtract (1/2) from both sides, we get
6m /7 +1/2 -1/2 = 1/2 -1/2
6m / 7 = 0
Multiply both sides by 7, we get
6m/7 * 7 = 0 * 7
6m = 0
Now divide the above equation by 6 we get,
6m / 6 = 0 /6
m = 0
Hence the value of "m" is 0